OpenStudy (shal):

1. Provide the reasons for the proof. Given: m angle1 = m angle2 Prove: Answer: Statements Reasons a. m angle 1 = m angle 2 a. Given b. angle 1 is supplementary to angle 2. b. Definition of supplementary angles. c. m angle 1 + m angle 2 = 180 c. d. m angle1 + m angle1 = 180 2(m angle1) = 180 d. e. m angle1 = 90 e. f. angle 1 is a right angle. f. GIVEN OR PROVE IF PROVE PUT REASON....EXAMPLE: SUPPLEMENTARY......

OpenStudy (shal):

FOR EACH LETTER

OpenStudy (shal):

THERE THTS THE LINK

OpenStudy (anonymous):

So, c, d, e, f, and g, right?

OpenStudy (shal):

NO IM NOT SURE OF B

OpenStudy (shal):

BUT A IS CORRECT I KNOE

OpenStudy (anonymous):

So, te first one you haven't proven is b's info right? The reason is that the lines intersect the same lines.

OpenStudy (anonymous):

Wait, are you given the statements, and needing to prove them?

OpenStudy (shal):

yes

OpenStudy (anonymous):

Okay, that makes life easier then. Do you see how I answered that?

OpenStudy (shal):

yes,

OpenStudy (anonymous):

SO, the next one, is the measures are equal. Can you use a and b to prove c?

OpenStudy (shal):

but can yu still help me with the other letters i have been waitin like for r2 hours in this website

OpenStudy (anonymous):

On a and b?

OpenStudy (shal):

so for which letter?

OpenStudy (anonymous):

I'm working step by step through the others

OpenStudy (shal):

oh ok so which one was meauring ?

OpenStudy (anonymous):

What do you mean?

OpenStudy (shal):

i mean which one is the one tht yu said for measuring

OpenStudy (shal):

i gave yu a medal

OpenStudy (shal):

lol

OpenStudy (anonymous):

Sorry, misread something, And thank you Statement c, is tgat measures of 1+2= 180*

OpenStudy (anonymous):

So, can you use either of the previous steps to prove this?

OpenStudy (shal):

yes

OpenStudy (anonymous):

Do it

OpenStudy (shal):

which one?

OpenStudy (shal):

which previous step?

OpenStudy (shal):

can yu fill up the doc and send it back

OpenStudy (shal):

so it could be easier

OpenStudy (shal):

becuz it gets confusing

OpenStudy (shal):

soryy

OpenStudy (anonymous):

Prove c, using the fact that the definition of supplementary angles means that the measures' sum equals 180* I will not, because that will not help you understand it. It's why we'redoing this slowly. If you have questions, just ask, and I'll address them best I can.

OpenStudy (shal):

ok, thx atleast yur helping me

OpenStudy (anonymous):

Does the answer to "prove c" make sense?

OpenStudy (shal):

yes

OpenStudy (anonymous):

Alright, so can you make a guess on how to do d, which is: d. m angle1 + m angle1 = 180 2(m angle1) = 180

OpenStudy (shal):

given

OpenStudy (anonymous):

Bingo! Because you know that m<1 = m<2, you can prove that 2(m<1)=180* if m angle1 + m angle1 = 180 by the def of supp angles

OpenStudy (shal):

cool, i wish yu couldbe my tutor

OpenStudy (anonymous):

It's really too bad there isn't a phone-like option, because typing all of this out is a little annoying

OpenStudy (shal):

yep

OpenStudy (shal):

but what can i do lol

OpenStudy (shal):

so e f g are left

OpenStudy (anonymous):

So, for e, you can figure that out right? I'll assume you can do simple math. F is the same, just using a definition

OpenStudy (shal):

i actually need help

OpenStudy (shal):

so f is proven?

OpenStudy (anonymous):

For g, the information is that any line intersecting another with at a right angle. Simply put, the definition of perpendicular lines, based on the fact that m<1=90=m<2

OpenStudy (anonymous):

As for f, it's simply that any angle =90* is a right angle, def of rt angles

OpenStudy (shal):

what was e?

OpenStudy (shal):

so e f g are all proven?

OpenStudy (anonymous):

E, is that: (2m<1= 180*)/2 m<1=90* You're really just dividing by two, or isolating the angle measure

OpenStudy (anonymous):

Yes they are

OpenStudy (shal):

ok

OpenStudy (shal):

Thansk so much!

OpenStudy (anonymous):

No problem, don't forget to close down this question, alright?

OpenStudy (shal):

ok and

OpenStudy (shal):

do yu think i will get a good grade?

OpenStudy (shal):

do yu have time to answer numbers 2 3 4 ?

OpenStudy (anonymous):

I think, if you slow down, and double check what you have, then you will do fine. But I'm hardly one to judge, as we've just met

OpenStudy (anonymous):

Sorry, I my own hwk to do

OpenStudy (shal):

ok thx!